3) Use the Jacobian to classify each of the equilibrium point(s). (spiral source, nodal sink, etc.) x = (1 + x) sin(y) y =

Please avoid the use of a plotter, but you can graph it by hand. Please explain how you got the equilibrium points in detail and thanks in advance! I've been stuck on this one for a few days.
Answer #1

у 1- (1 +1 ) sin(1) P(uy 8 = 1-4. — єфу eos (1 ) = 4 () Far fixed point plavy) =0 44 см, )=0 (1+1) *** *> siny yant . К k qwthe critical points are for nevenas (0,07) & for nodd (2,07) Now Jony Pre Py Que Qy siny (1+) costs -1 sing conn) For n even